OpenStudy (anonymous):

let f(x,y)= (x+y)^2/ x-y we claim that it cannot be extended over (0,0), is that true?

OpenStudy (perl):

What do you mean by 'cannot be extended over (0,0)' the limit might exist, so you could redefine f(0,0)

OpenStudy (anonymous):

f(0,0) is undefined but if the claim true i need help finding a curve or direction from our domain which is |x|>|y|

OpenStudy (anonymous):

what can we choose y from the domain to show that the claim is true?

OpenStudy (perl):

Sorry I'm having trouble understanding the language of the claim.

OpenStudy (perl):

Are you trying to extend the function so it is continuous?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for example let y= mx where 0<=m<1 so the lim will go to 0

OpenStudy (anonymous):

what else can we choose y to be?

OpenStudy (perl):

Ok, let's try that. y = x

OpenStudy (anonymous):

but our domain is |x|>|y|, is that work?

OpenStudy (perl):

if y = x we have (2y)^2 / (y - y ) = 2y^2 / 0 , which is a nonremovable discontinuity

OpenStudy (perl):

I don't think it should make a difference that |x| > |y| .

OpenStudy (anonymous):

so that is not continuous so it cannot extend over (0,0)?

OpenStudy (anonymous):

what if we let y= x/2 , does it work?

OpenStudy (across):

Sure. Just let f(x,y)=(0,0) at (0,0).

OpenStudy (anonymous):

I did that but my professor asked for a curve or direction from our domain which is |x|>|y| that satisfies the claim

OpenStudy (across):

sorry. i absolutely misread your question. have you considered \[\gamma\left(t\right)=\left(x,x\cos x\right)?\] It's a smooth curve such that\[\lim_{t\to0^-}f\left(\gamma\left(t\right)\right)=-\infty\qquad\text{and}\qquad\lim_{t\to0^+}f\left(\gamma\left(t\right)\right)=+\infty.\]

OpenStudy (across):

This shows that \(f\) cannot be extended at \((0,0)\)